# Non-Linear Flow Matching for Full-Atom Peptide Design

### v1

@@ 3.3 Linear Flow for Orientation and Amino Acid Type
Amino Acid Types.
@@
3.4 Loss Balancing for Sequence-Structure Co-design
@@
3.5 Sampling with ODE
We perform sampling with the pre-trained flow model by formulating peptide generation as an ODE, where the peptide ${Lig}$ evolves according to $\frac{d}{dt}Lig=v_{t}({Lig})$, with derivatives computed for each of the four modalities. The equation is discretized into $N$ steps, yielding the final peptide ¯$\overline{}{{Lig}_{1}}$.
Each modality’s update depends not only on its own state but also on the states of other modalities. This interdependence highlights the importance of accurately capturing the temporal relationships across modalities. By iteratively updating these states, our model effectively simulates the dynamic nature of peptide docking, while enabling flexible sampling strategies for different design tasks (fix sequence for re-docking and fix backbone for side-chain packing).
@@ 3.5 Sampling with ODE
  Init: Initial state $Lig_{0}=\{(X_{i}^{0},R_{i}^{0},\boldsymbol{\tau}_{i}^{0},a_{i}^{0})\}_{i=1}^{n}$
  for $t=1$ to $N$ do
     Predict gradient vecto $v_{t}=\text{Dec}(Lig_{\frac{t-1}{N}},Rec,t)$
     EulerStep ${Lig_{\frac{t}{N}}}=Lig_{\frac{t-1}{N}}+\Delta t\cdot v$
@@
  Output: $\overline{{Lig}_{1}}$
Table 1: Comparison for Sequence-Structure Co-design: Highest Affinity and Stability.
Method Stability $\uparrow$ Affinity $\uparrow$ Novelty $\uparrow$ Diversity Seq sim
@@
NLFlow 12.33% 17.81% 84.94% 0.651 44.61%
NLFlow w/o II 10.42% 11.11% 85.34% 0.648 48.26%
3.6 Network Parametrization
Encode with Interaction-related Information.
We use two multi-layer perceptrons (MLPs) to encode the features of amino acids and the relationships between residue pairs. The first MLP encodes the features of individual residues, processing the amino acid type, backbone dihedral angles, and local atomic coordinates. Position embeddings are integrated to identify the context of sequences. Additionally, properties such as polarity, charge, hydrophilicity, and sulfur presence (which is critical for stability) are encoded using one-hot vectors, serving as auxiliary interaction-related information. This allows the model to not only focus on structural alignment but also on the alignment of interaction forces, ultimately enhancing the affinity between the peptide and protein. The second MLP encodes the relationships between residue pairs, capturing their relative positions, distances between atoms, and dihedral angles.
Learning Gradient Vector with IPA and Transformers.
We use invariant point attention (IPA) (Lee et al., 2019b) and transformer to learn the gradient vector field $v_{t}$. The model consists of multiple IPA blocks for spatial feature learning, followed by transformer encoders for sequence modeling. Each block includes layer normalization, transition layers, and backbone updates, iteratively refining peptide conformations. Residue identity, angular encoding, and time-step embedding are integrated to ensure smooth temporal updates.
4 Experiment
4.1 Set up
In this section, we evaluate NLFlow on four tasks: (i) sequence-structure co-design, (ii) re-docking, (iii) side-chain packing and (iv) one-step generation. Through these evaluations, we assess whether the correct temporal characterization leads to the generation of more stable, high-affinity peptides with accurate structural alignment and functional relevance.
The dataset, derived from the work of PepFlow (Li et al., 2024), was obtained from PepBDB (Wen et al., 2019) and Q-BioLip (Wei et al., 2024). To ensure high-quality data, duplicates were removed, a resolution threshold of less than 4 Åwas applied, and peptide lengths were restricted to between 3 and 25 residues. This preprocessing resulted in a final dataset consisting of 10,348 complexes, with 166 complexes reserved for the test set, and the remaining data split for training and validation.
During pre-training, we simply selected the hyperparameter $k=2$ for the non-linear flow model discussed in section 3.2. We trained three variants of our model to evaluate different configurations: (i) NLFlow, the full model with all components. (ii) NLFlow w/o II, which excludes the interaction-force related information. (iii) NLFlow w/o II+LW, which further removes the sequence-structure balance weight. Each model variant was trained for a total of 65,000 iterations to ensure sufficient training and convergence. The experiments were conducted on a Tesla P40 GPU with a batch size of 12, using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a plateau learning rate scheduler with a factor of 0.8, patience of 10, and a minimum learning rate of $5\times 10^{-6}$.
@@ 4.1 Set up
Table 2: Comparison for Peptide Re-docking. Suc. $=$Success, Div. $=$ Diversity
@@
NLFlow 4.09 80.01% 85.48% 52.15% 0.638
NLF w/o II 4.56 80.64% 85.67% 60.12% 0.647
NLF w/o II+LW 4.41 80.91% 85.63% 58.89% 0.624
Table 3: Comparison for Side-chain Packing
@@
NLFlow 11.69 19.66 54.87 55.74
4.2 Sequence-Structure Co-design
Sequence-structure co-design involves jointly generating the peptide sequence and conformation, resulting in a full-atom peptide docked onto the target protein. We evaluate three baseline models: DiffPP, PPFlow (Lin et al., 2024) and PepFlow (Li et al., 2024). DiffPP is a diffusion model for protein backbone parametrization, using DDPM (Yang et al., 2023) and SO(3)-DPM (Leach et al., 2022) to model translation and rotation, along with multinomial diffusion for amino acid types. PepFlow and PPFlow, which were proposed simultaneously, are the latest models based on the flow matching framework for peptide design. They represent peptide structures by modeling backbone frames on the SE(3) manifold and side-chain dynamics on high-dimensional tori, enabling the generation of full-atom peptides with a focus on structural accuracy and torsion angle optimization.
Metrics.
The evaluation of the generated peptides is based on five metrics. Energy is calculated using Rosetta (Rohl et al., 2004), with two primary measures: Affinity, which quantifies the binding energy, where lower values indicate stronger binding potential, and Stability, representing the overall energy of the peptide-protein complex, with lower values indicating a more stable complex. We report the percentages of designed peptides with higher affinity and stability than the reference ones. Diversity is evaluated using the TM-score (Zhang & Skolnick, 2005), defined as the average of 1 minus the pairwise TM-scores of generated peptides, with Novelty calculated as the percentage of TM-score less than 0.5. Seq Sim is quantified by the longest common subsequence ratio between peptide pairs, reflecting the consistency between sequence and conformation.
Results: Achieves Superior Affinity and Stability.
From the comparison in Table 1, it can be concluded that (i) NLFlow exhibits a significant advantage in energy metrics, achieving 6.99% higher stability and 4.62% higher affinity compared to baseline models. A key distinction from flow-based methods explains that this improvement is largely due to the effective handling of temporal inconsistencies. (ii) The inclusion of interaction-related information (II) results in the generation of 6.7% more peptides with lower binding affinity compared to the version w/o II, highlighting its contribution to improved peptide-protein interactions. (iii) Our method generates diverse and novel peptide conformations. In contrast, the stronger randomness of diffusion models enables broader exploration of the conformational space, while NLFlow focuses on finer adjustments near the pocket, leading to slightly lower diversity. (iv) The designed sequences also show corresponding differences, reflecting the consistency between sequence and structural in the co-design process.
4.3 Peptide Re-docking
The re-docking task evaluates the model’s ability to reconstruct the conformation by fixing the reference peptide sequence during the sampling stage, without retraining the model, and generating a full-atom peptide-protein complex in the docked state. We evaluate two baseline models in re-docking task: PepFlow and HDOCK (Yan et al., 2020). HDOCK is a traditional docking method, which uses a combination of rigid-body docking followed by energy minimization to predict the binding mode of two interacting proteins.
Metrics.
The evaluation of the Peptide Re-docking task is based on five metrics: RMSD measures the structural deviation between the generated peptide’s $C_{\alpha}$ atoms and the reference peptide, indicating conformation accuracy; SSR quantifies the similarity in secondary structure, assessing the preservation of secondary structural features; BSR calculates the overlap between the docking site of the generated and the reference peptide-protein complex, reflecting binding site accuracy; Success is defined as achieving a top-1 RMSD $<$ 4 Å, with BSR and SSR both greater than 0.8. The result reports the percentage of successful cases; and Diversity, same as the co-design task, calculated as 1 minus the average TM-score, measuring the variation in peptide conformations.
Results: Lower RMSD in Reconstructing.
Table 3 shows that (i) NLFlow achieves the lowest RMSD, performing superiorly in structural accuracy. This demonstrates that by addressing temporal inconsistencies between modalities, the correct docking process simulation provides the model with a strong ability to reconstruct conformations. As illustrated in Figure 4, NLFlow not only focuses on fine-grained atomic coordinates but also effectively restores the overall macrostructural features, reflecting the model’s deep understanding of the docking task under correct temporal alignment. While PepFlow excels in binding site recognition (BSR), NLFlow achieves a better balance between structural alignment and conformational diversity. (ii) Compared to flow-based models, the traditional method HDOCK, which does not redesign the peptide structure, does not provide SSR or diversity results, limiting its ability to model peptide-protein interactions. (iii) Among our model variants, removing interaction force information (w/o II) increases the success rate, as it shifts the model’s focus from balancing structural alignment and docking interactions to solely optimizing structural alignment.
4.4 Side-chain Packing
This task evaluates the model’s ability to predict the correct torsional angles for the side chains, which are crucial for accurate protein-ligand docking and stability. Specifically, we calculate the mean squared error (MSE) of the four predicted side-chain torsional angles. We use energy-based methods: RosettaPacker (Leman et al., 2020), SCWRL4 (Krivov et al., 2009), and Rotamer Density Estimator (RDE) (Luo et al., 2023) with Conditional Flow on TNrt: RDE-PP (Lin et al., 2024) as baselines.
Results: Strong Performance in Lower-Order Torsions.
@@ 4.4 Side-chain Packing
As shown in Table 3, NLFlow demonstrates a clear advantage in predicting the $\chi_{1}$ and $\chi_{2}$ torsional angles, outperforming all other methods. This highlights the importance of adjusting torsion angles according to the pocket environment, reflecting NLFlow’s ability to finely tune the conformation near the pocket. However, for the more challenging $\chi_{3}$ and $\chi_{4}$ angles, the performance of all models is comparable, with showing worse results compared to $\chi_{1}$ and $\chi_{2}$. This highlights that higher-order torsions remain a complex task, with no model consistently outperforming others across all angles.
4.5 One-step Generation
@@
Most flow matching methods rely on the assumption that the direction of the gradient vector field remaining constant, which theoretically enables one-step generation. However, achieving this in practice is challenging in non-linear flow due to two key issues: first, path conflicts as identified by Rectified Flow (Liu et al., 2022) exist across all flow-based models; second, our non-linear assumptions hinder the ability of the initial time step to capture temporal inconsistencies across different modalities, limiting the effectiveness of one-step generation.
@@
Results: Rapid Improvement with Minimal Steps.
As shown in Fig 5, at $step=1$, the lack of temporal consistency across modalities leads to poor performance, particularly with stability (Stab = 0). However, by $step=2$, the model rapidly improves, indicating that the difficulty of one-step generation primarily arises from the absence of temporal information, highlighting the importance of the non-linear assumption. After $step>10$, further increasing the number of steps has minimal impact on the model’s performance, suggesting that our model largely follows the straight-line assumption of flow, making it a fast and efficient generation model.
5 Conclusion
In this work, we present NLFlow, a non-linear flow matching framework designed to address the critical challenge of temporal inconsistency across modalities in full-atom peptide design. By introducing a polynomial-based interpolation scheme and its associated time-varying gradient vector field, NLFlow explicitly decouples the evolution of the positional manifold from rotational, torsional, and residue-type manifolds. This approach captures the biologically hierarchical nature of peptide docking, resulting in peptides with enhanced binding affinity and stability. The integration of interaction-related features further enables the model to balance structural alignment with functional optimization, providing a nuanced simulation of peptide-protein interactions.
Despite these advancements, NLFlow remains limited by suboptimal accuracy in predicting higher-order torsions under complex conformational constraints, and its predefined polynomial interpolation, though effective for temporal decoupling, restricts flexibility in modeling dynamic temporal hierarchies. Future directions include developing adaptive flow architectures to learn temporal dynamics directly from data, and incorporating dynamic interaction force predictors for context-aware refinement.
Appendix A Examples of peptides
We present some other peptide results generated in the re-docking task, showing the recovery of the structure.
Figure 6: Examples of peptides generated in the re-docking task

### v2

@@ III-C2 Torsion Angles
III-C3 Amino Acid Types
@@
III-D Loss Balancing for Sequence-Structure Co-design
@@
III-E Sampling with ODE
We perform sampling with the pre-trained flow model by formulating peptide generation as an ODE, where the peptide Lig{Lig} evolves according to $\frac{d}{dt}Lig=v_{t}({Lig})$, with derivatives computed for each of the four modalities. The equation is discretized into $N$ steps, yielding the final peptide $\overline{}{{Lig}_{1}}$.
Each modality’s update depends not only on its own state but also on the states of other modalities. This interdependence highlights the importance of accurately capturing the temporal relationships across modalities. By iteratively updating these states, our model effectively simulates the dynamic nature of peptide docking, while enabling flexible sampling strategies for different design tasks such as fix sequence for re-docking, and fix backbone for side-chain packing.
@@ III-E Sampling with ODE
Init: Initial state $Lig^{0}=\{(X_{i}^{0},R_{i}^{0},\boldsymbol{\tau}_{i}^{0},a_{i}^{0})\}_{i=1}^{n}$
for $t=1$ to NN do
  Predict gradient vecto $v_{t}=\text{Dec}(Lig^{\frac{t-1}{N}},Rec,t)$
  EulerStep ${Lig^{\frac{t}{N}}}=Lig^{\frac{t-1}{N}}+\Delta t\cdot v$
@@
Output: $\overline{{Lig}^{1}}$
III-F Network Parametrization
III-F1 Encode with Interaction-related Information
We use two multi-layer perceptrons (MLPs) to encode the features of amino acids and the relationships between residue pairs. The first MLP encodes the features of individual residues, processing the amino acid type, backbone dihedral angles, and local atomic coordinates. Position embeddings are integrated to identify the context of sequences. Additionally, properties such as polarity, charge, hydrophilicity, and sulfur presence, which is critical for stability, are encoded using one-hot vectors, serving as auxiliary interaction-related information. This allows the model to not only focus on structural alignment but also on the alignment of interaction forces, ultimately enhancing the affinity between the peptide and protein. The second MLP encodes the relationships between residue pairs, capturing their relative positions, distances between atoms, and dihedral angles.
III-F2 Learning Gradient Vector with IPA and Transformers
We use invariant point attention (IPA) [28] and transformer to learn the gradient vector field $v_{t}$. The model consists of multiple IPA blocks for spatial feature learning, followed by transformer encoders for sequence modeling. Each block includes layer normalization, transition layers, and backbone updates, iteratively refining peptide conformations. Residue identity, angular encoding, and time-step embedding are integrated to ensure smooth temporal updates.
IV Experiment
IV-A Set up
In this section, we evaluate THFlow on four tasks: (i) sequence-structure co-design, (ii) re-docking, (iii) side-chain packing and (iv) one-step generation. Through these evaluations, we assess whether the correct temporal characterization leads to the generation of more stable, high-affinity peptides with accurate structural alignment and functional relevance.
The dataset, derived from the work of PepFlow [10], was obtained from PepBDB [29] and Q-BioLip [30]. To ensure high-quality data, duplicates were removed, a resolution threshold of less than 4 Åwas applied, and peptide lengths were restricted to between 3 and 25 residues. This preprocessing resulted in a final dataset consisting of 10,348 complexes, with 166 complexes reserved for the test set, and the remaining data split for training and validation.
During pre-training, we simply selected the hyperparameter $k=2$ for the time-varying flow model discussed in section III-B. We trained three variants of our model to evaluate different configurations: (i) THFlow, the full model with all components. (ii) THFlow w/o II, which excludes the interaction-force related information. (iii) THFlow w/o II+LW, which further removes the sequence-structure balance weight. Each model variant was trained for a total of 65,000 iterations to ensure sufficient training and convergence. The experiments were conducted on a Tesla P40 GPU with a batch size of 12, using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a plateau learning rate scheduler with a factor of 0.8, patience of 10, and a minimum learning rate of $5\times 10^{-6}$.
IV-B Sequence-Structure Co-design
Sequence-structure co-design involves jointly generating the peptide sequence and conformation, resulting in a full-atom peptide docked onto the target protein. We evaluate three baseline models: DiffPP, PPFlow [11] and PepFlow [10]. DiffPP is a diffusion model for protein backbone parametrization, using DDPM [31] and SO(3)-DPM [32] to model translation and rotation, along with multinomial diffusion for amino acid types. PepFlow and PPFlow, which were proposed simultaneously, are the latest models based on the flow matching framework for peptide design. They represent peptide structures by modeling backbone frames on the SE(3) manifold and side-chain dynamics on high-dimensional tori, enabling the generation of full-atom peptides with a focus on structural accuracy and torsion angle optimization.
TABLE I: Comparison for Sequence-Structure Co-design: Highest Affinity and Stability. Stab. == Stability, Aff. == Affinity, Nov. == Novelty, Div. == Diversity
Method Stab. $\uparrow$ Aff. $\uparrow$ Nov. $\uparrow$ Div. AAS
@@ IV-B Sequence-Structure Co-design
THFlow 12.33% 17.81% 84.94% 0.651 44.61%
THF w/o II 10.42% 11.11% 85.34% 0.648 48.26%
THF w/o II+LW 4.32% 7.19% 86.23% 0.645 49.27%
IV-B1 Metrics
The evaluation of the generated peptides is based on five metrics. Energy is calculated using Rosetta [33], with two primary measures: Affinity, which quantifies the binding energy $\Delta G$, where lower values indicate stronger binding potential, and Stability, representing the overall energy of the peptide-protein complex, with lower values indicating a more stable complex. We report the percentages of designed peptides with higher affinity and stability than the reference ones. Diversity is evaluated using the TM-score [34], defined as the average of 1 minus the pairwise TM-scores of generated peptides, with Novelty calculated as the percentage of TM-score less than 0.5. AAS refers to the similarity of amino acid sequences between generated and ground truth peptide pairs, quantified as the longest common subsequence ratio. Co-analysis with structural novelty can reflect the consistency between sequence and conformation.
IV-B2 Results: Achieves Superior Affinity and Stability
From the comparison in Table I, it can be concluded that (i) THFlow exhibits a significant advantage in energy metrics, achieving 6.99% higher stability and 4.62% higher affinity compared to baseline models. A key distinction from flow-based methods explains that this improvement is largely due to the effective handling of multimodal temporal inconsistency problem. (ii) Our method generates diverse and novel peptide conformations. In contrast, the stronger randomness of diffusion models enables broader exploration of the conformational space, while THFlow focuses on finer adjustments near the pocket, leading to slightly lower diversity. (iii) The designed sequences also show corresponding differences, reflecting the consistency between sequence and structural in the co-design process. (iv) Ablation experiments show the positive effects of interaction-related information and sequence-structure balance weight modules on the model. The former can deepen the understanding of the physical and chemical significance of the binding interface, and bring the improvement of the affinity. The latter shows the importance of the interaction between structure and sequence in training. After removing the constraints, the sequence generated by the model is more similar to the ground truth, but the structure is more novel, indicating that there are many unreasonable new structures generated in the same environment, which is reflected in the worse affinity and stability.
IV-C Peptide Re-docking
The re-docking task evaluates the model’s ability to reconstruct the conformation by fixing the reference peptide sequence during the sampling stage, without retraining the model, and generating a full-atom peptide-protein complex in the docked state. We evaluate two baseline models in re-docking task: PepFlow and HDOCK [35]. HDOCK is a traditional docking method, which uses a combination of rigid-body docking followed by energy minimization to predict the binding mode of two interacting proteins.
@@ IV-C Peptide Re-docking
TABLE II: Comparison for Peptide Re-docking: Lowest RMSD. Suc. ==Success, Div. == Diversity
@@
THFlow 4.09 80.01% 85.48% 52.15% 0.638
THF w/o II 4.56 80.64% 85.67% 60.12% 0.647
THF w/o II+LW 4.41 80.91% 85.63% 58.89% 0.624
IV-C1 Metrics
The evaluation of the peptide re-docking task is based on five metrics: RMSD measures the structural deviation between the generated peptide’s $C_{\alpha}$ atoms and the reference peptide, indicating conformation accuracy; SSR quantifies the similarity in secondary structure, assessing the preservation of secondary structural features; BSR calculates the overlap between the docking site of the generated and the reference peptide-protein complex, reflecting binding site accuracy; Success is defined as achieving a top-1 RMSD << 4 Å, with BSR and SSR both greater than 0.8. The result reports the percentage of successful cases; and Diversity, same as the co-design task, calculated as 1 minus the average TM-score, measuring the variation in peptide conformations.
IV-C2 Results: Lower RMSD in Reconstructing
Table II shows that (i) THFlow achieves the lowest RMSD, this demonstrates that the correct docking process simulation provides the model with a strong ability to reconstruct conformations. (ii) As illustrated in the right of Figure 4, we found many examples showing THFlow’s attempts to understand and reduce the secondary structure of peptides. Although this may result in a higher RMSD, we consider it as the model exploring the association between coordinates, structure, and interactions. This also explains why our model has higher diversity. (iii) Removing interaction information (w/o II) improves the success rate, likely because it eliminates a difficult trade-off. This allows the model to dedicate its full capacity to optimizing structural alignment instead of balancing it with docking forces.
IV-D Side-chain Packing
This task evaluates the model’s ability to predict the correct torsional angles for the side chains, which are crucial for accurate protein-ligand docking and stability. Specifically, we calculate the mean squared error (MSE) of the four predicted side-chain torsional angles. We use energy-based methods: RosettaPacker [36], SCWRL4 [37], and Rotamer Density Estimator (RDE) [17] with Conditional Flow on TNrt: RDE-PP [11] as baselines.
TABLE III: Comparison for Side-chain Packing
@@ IV-D Side-chain Packing
THFlow 11.69 19.66 54.87 55.74
IV-D1 Results: Strong Performance in Lower-Order Torsions
@@
As shown in Table III, THFlow demonstrates a clear advantage in predicting the $\chi_{1}$ and $\chi_{2}$ torsional angles, outperforming all other methods. This highlights the importance of adjusting torsion angles according to the pocket environment, reflecting THFlow’s ability to finely tune the conformation near the pocket. However, for the more challenging $\chi_{3}$ and $\chi_{4}$ angles, the performance of all models is comparable, with showing worse results compared to $\chi_{1}$ and $\chi_{2}$. This highlights that higher-order torsions remain a complex task, with no model consistently outperforming others across all angles.
IV-E One-step Generation
@@
Most flow matching methods rely on the assumption that the direction of the gradient vector field remaining constant, which theoretically enables one-step generation. However, achieving this in practice is challenging due to two key issues: first, path conflicts as identified by Rectified Flow [38] exist across all flow-based models; second, our polynomial assumptions hinder the ability of the initial time step to capture temporal inconsistencies across different modalities, limiting the effectiveness of one-step generation.
@@
IV-E1 Results: Rapid Improvement with Minimal Steps
As shown in Fig 5, at $step=1$, the lack of temporal consistency across modalities leads to poor performance, particularly with stability (Stab = 0). However, by $step=2$, the model rapidly improves, indicating that the difficulty of one-step generation primarily arises from the absence of temporal information, highlighting the importance of the temporally hierarchical assumption. After $step>10$, further increasing the number of steps has minimal impact on the model’s performance, suggesting that our model largely follows the straight-line assumption of flow, making it a fast and efficient generation model.
V Conclusion
In this work, we present THFlow, a temporally hierarchical flow matching framework designed to address the critical challenge of temporal inconsistency across modalities in full-atom peptide design. By introducing a polynomial-based interpolation scheme and its associated time-varying gradient vector field, THFlow explicitly decouples the evolution of the positional manifold from rotational, torsional, and residue-type manifolds. This approach captures the biologically hierarchical nature of peptide docking, resulting in peptides with enhanced binding affinity and stability. The integration of interaction-related features further enables the model to balance structural alignment with functional optimization, providing a nuanced simulation of peptide-protein interactions.
Despite these advancements, THFlow remains limited by suboptimal accuracy in predicting higher-order torsions under complex conformational constraints, and its predefined polynomial interpolation, though effective for temporal decoupling, restricts flexibility in modeling dynamic temporal hierarchies. Future directions include developing adaptive flow architectures to learn temporal dynamics directly from data, and incorporating dynamic interaction force predictors for context-aware refinement.